The EMF of the cell
Cd ( CdCl2. 2.1/2 H2O (saturated) AgCl (s) ( Ag


is found to be 0.6753 V at 250C and 0.6915 V at 00C. Calculate (G, (H and (S of the cell reaction at 250C.
1
Expert's answer
2020-05-08T14:10:54-0400

Cell reaction is:

Cd + 2 AgCl = CdCl2 + 2 Ag

EMF corresponds to ΔG:

ΔG = -2*F*EMF = ΔH - T*ΔS

At 25 0C = 298 K we have the one equation: -2*96500*0.6753 = ΔH - 298*ΔS

At 0 0C = 273 K we have the other equation: -2*96500*0.6915 = ΔH - 273*ΔS

Solving the system, we have that at 25 0C = 298 K:

ΔH = -167601.972 J

ΔS = -125.064 J/K

ΔG = -2*96500*0.6753 = -130332.9 J